OpenStudy (anonymous):

what is the slope of the line 9x - 4 lny =3 that goes through the pt (1/3, 1) ?

OpenStudy (anonymous):

ax + by + c = 0 Slope = -a / b

OpenStudy (shubhamsrg):

@Yahoo! you misunderstood, it lny there, so its a curve, we are asked to find the slope of tangent at (1/3 ,1) for that express the eqn in terms of y (9x-3)/4 = lny =>y = e^(9x-3)/4 not find dy/dx in that put x=1/3 and y=1

OpenStudy (shubhamsrg):

now *

OpenStudy (anonymous):

o.O....Sorry